OpenStudy (anonymous):

The midpoint of a line segment with end points as (-1, 5) and (x2, 7) is (-4, 6). What is the value of x2? The midpoint of a line segment with end points as (-1, 5) and (x2, 7) is (-4, 6). What is the value of x2? @Mathematics

OpenStudy (anonymous):

(-1+x2)/2=-4 -1+x2=-8 x2=-7

OpenStudy (anonymous):

thank you! (: